OpenStudy (anonymous):

A university bookstore recently sold a wirebound graph-paper notebook for $2.15, and a college-ruled notebook for $3.43. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $134.38. How many of each type were sold?

OpenStudy (anonymous):

i want to know if my work write

OpenStudy (anonymous):

W+C=50 2.15W+3.43C=134.38 C=50-W 2.15W+3.43(50-W)=134.38 2.15W-3.43W=134.38-171.5 -1.28W=-37.12 W=29 C=50-29=21 W=29 C=21

OpenStudy (amistre64):

w + c = t wx + cy = T ; eliminate one of the variables -wx -cx = -tx wx +cy = T ------------- c(y-x) = T-tx c = (T-tx)/(y-x) c = (134.38 - 50(2.15))/(3.43-2.15) = 21 your right

OpenStudy (anonymous):

ty